The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 34 in., find the dimensions of the rectangle.
Answer by ptaylor(2198)

Let x=width
Then 2x+2=Length
Perimeter=34=2L+2W So our equation to solve is:
34=2(2x+2)+2(x) get rid of parens
34=4x+4+2x subtract 4 from both sides
34-4=4x+4-4+2x collect like terms
30=6x divide both sides by 6
x=5 in----------------------------------width
2x+2=2(5)+2=12 in--------------------------length
ck
2(5)+2(12)=34
10+24=34
34=34

Hope this helps-----ptaylor